If x and y are arbitrary real numbers with x < y, prove that there exists at least one irrational number z satisfying x < z < y, and hence infinitely many.


1
Expert's answer
2021-01-20T03:20:08-0500

Statement 1 (Density of rational numbers in "\R" )

If "u\in\R" and "v\in \R" with "u<v," then there exists at least one rational number "w" satisfying "u<w<v," and hence infinitely many.


Statement 2

"\sqrt{2}" is not a rational number. In general, if "p" is a prime number, then "\sqrt{p}" is not a rational number.


Statement 3

If "a" is an arbitrary rational number and "b" is an arbitrary irrational number, then "ab" is irrational provided "a\not=0."


If we consider "u=x/\sqrt{2}" and "v=y/\sqrt{2}" then it follows from Statement 1 that there exists rational number "w\not=0" such that 


"x/\sqrt{2}<w<y/\sqrt{2}"

Then


"x<w\sqrt{2}<y"

If we set "z=w\sqrt{2}" then it follows from Statement 2 and Statement 3 that "z" is irrational number.


Therefore

If "x\in \R" and "y\in\R" with "x<y," then there exists at least one irrational number "z" satisfying "x<z<y," and hence infinitely many.